A current of 75.8 mA exists in a metal wire.

(a) How many electrons flow past a given cross section of the wire in 10.8 min?
electrons

(b) In what direction do the electrons travel with respect to the current?

same directionopposite direction    The magnitude is zero.
Answer #1

Current is given by charge flow per unit time. Here the charge is electron so it is electron flow per unit time.

I = Q / t

so Q = I * t = 0.0758 A * 10.8 * 60 sec

Q = 49.1184 C

Part (B)

electron always flow in the negative direction of the current so here answer is " opposite direction "

feel free to drop comment if you have any doubt.